OpenStudy (anonymous):

Solve the following trinomial: x2 + 5x + 6 = 0

OpenStudy (mimi_x3):

(x+3)(x+2) x = - 3 , -2

OpenStudy (anonymous):

(x+2) (x+3) x= -2 and x= -3

OpenStudy (anonymous):

\[x^2+5x+6\] \[x^2+3x+2x+6\] \[x(x+3)+2(x+3)=0\] \[(x+2)(x+3)=0\] either x+2 is 0 or x+3 is 0 \[x+2=0, x=-2\] \[x+3=0 , x=-3\]